Fermi National Laboratory


Accelerator Update - July 21-July 23

The integrated luminosity for last week was 5106.9 nb-1.

Monday July 21
The day shift began with the Antiproton Source (Pbar) stacking, with Tevatron (TeV) experts conducting studies, and with MiniBooNE taking beam.

The TeV quenched at 8:46 AM while operations was trying to terminate a 36x0 study store. The quench occurred at TeV sector BA.

At 9:32 AM, cryo experts installed new software for controlling cryo valves that should stop the occasional valve excursions that have caused quenches in the past. If this test code works on the sector A cryo system, it will be installed everywhere.

Both D-zero (D0) and the Collider Detector Facility (CDF) experiments had accessed their collision hall around 9:42 AM. Both experiments were out of access by 11:30 AM.

Cryo experts began conducting wet engine repairs around the TeV ring at 10:23 AM.

Operations put the TeV into shot setup and at 3:27 PM and established store 2810 with an initial luminosity of 31.65E30. Pbar resumed stacking.

At 9:13 PM, the Pbar stack went away. Operators investigated and discovered an interlock problem. Operators called in a safety system technician who replaced a failed summation switch at 11:30 PM.

Tuesday July 22
The midnight shift began with Operations monitoring store 2810, with MiniBooNE taking beam, and with operators interlocking the Pbar tunnel.

Cryo techs at CHL noticed a low speed alarm occasionally showing on the sector B3 wet engine. This will have to be repaired later.

Booster suffered a vacuum burst on RF stations 13 and 14 at 1:44 PM. Everything reset okay.

Operations conducted a quick end-of-store study and then terminated store 2810 at 2:09 PM.

CDF and D0 accessed their collision halls at 2:26 PM. Both experiments came out of access about thirty minutes later.

At 3:03 PM, a Main Injector RF station (MIRF9) started tripping off. Operations asked an expert to investigate.

Operations put the TeV through a dry squeeze at 3:40 PM.

The Linac was having a problem with its beam drifting at 5:26 PM. An expert began tuning the beam line.

Booster RF stations 5 and 6 tripped off due to a high vacuum alarm. Everything reset okay.

Operations established store 2813 at 7:40 PM with an initial luminosity of 31.64E30. Pbar resumed stacking.

Wednesday July 23
The midnight shift began with Operations monitoring store 2813, with Pbar stacking, and with MiniBooNE taking beam.

A Linac RF station (KRF4) tripped off at 7:27 AM. Operators responded and beat up an offending LCW flow switch. The station was off for about ten minutes.

Switchyard experts were preparing to conduct beam studies at 8 AM.

Today's Plan
The plan for today is to terminate store 2813 at 4 PM and then put the TeV into shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/23/2003   email Fermilab